A significantly over-exposed radiograph is produced using 300 mA, 0.20 seconds at 100 kVp. What new exposure time would be necessary to reduce IR exposure to ¼ the original, if the radiographer was required to employ 200 mA at 115 kVp.
Answer #1

Pfinal=1/4 Pinitial (P= Power)

200*115 / t = 300 * 100 /0.2

t= 0.1533 seconds